OpenStudy (anonymous):

A feather is dropped onto the surface of the moon. How far will the feather have fallen if it reaches the surface in 9.00 seconds? Given: g on moon = -1.63 meters/second2

OpenStudy (carmuz):

Take in count that in the moon there is no air resistance. The feather will fall as it were a hammer (or any other heavy object). Now: \[h = V_0t + \frac{ 1 }{ 2 } g_{moon}t^2\] The feather is dropped, so Vo = 0. \[h = \frac{ 1 }{ 2 } g_{moon}t^2\] \[h = \frac{ 1 }{ 2 } (1.63 m/s^2) (9s)^2\] \[h = 66 m\]

OpenStudy (carmuz):

:)